OpenStudy (anonymous):

Find the value of X: 35=(3-2)(3x) 3 -3 7 -7 (5 and 2 are high)

OpenStudy (igreen):

Oh, so what's 3^-2?

OpenStudy (mathmate):

I think you mean (5, -2 and x are high): \(\large 3^5=3^{-2} 3^x\)

OpenStudy (anonymous):

Idk @iGreen and yes @mathmate

OpenStudy (anonymous):

and its 3-2 not negative 2. Can you help?

OpenStudy (mathmate):

so it is \(\large 3^5=(3−2)3^x\)?

OpenStudy (anonymous):

I think? thats why i asked Lol. I have to find the value of X

OpenStudy (butterflydreamer):